# VHE gamma-ray intranight variability from BL Lacertae during the extreme flaring state of 2022

### 25 Mar 2026 v2

@@ Abstract
Context. BL Lacertae (BL Lac), the archetype blazar of its subclass and one of the most studied blazars in the last decades, has gone through a series of major multi-wavelength outbursts since 2020, resulting in its highest recorded -ray flare up to date between September and November 2022 together with those from August 2021 and October 2024.
Aims. We characterise the -ray and multi-wavelength emission and spectral energy distribution (SED) of BL Lac, as well as their evolution during the major and extended -ray and multi-wavelength flare occurring between September and November 2022.
Methods. We evaluate the variability of the flare, with focus on the nights of October 20 and November 13, when clear intranight very-high-energy (VHE, GeV) -ray variability is observed. We model the -ray and broadband SEDs during periods of stable emission identified with a Bayesian block analysis, interpreting their evolution of the flare from the variability of the relativistic particles and physical parameters of the jet.
Results. The VHE emission shows an average flux of 0.23 Crab Units (C.U.) above 200 GeV during this flare and a variability amplitude of more than a factor 10. Intranight doubling-flux variations as fast as 8 minutes are observed during the nights of October 20 and November 13, when maximum fluxes of 4.4 C.U. above 100 GeV and 2.8 C.U. above 200 GeV are reached. The spectral analysis reveals a transition of the X-ray emission from the high- to the low-energy SED peak, and a shift of the -ray peak towards higher energies. The broadband emission was interpreted within a leptonic two-zone model in which intranight variability is explained as magnetic reconnection in a compact region closely oriented with the line of sight, while weekly-scale variations can be explained as variations of the relativistic electron distributions and the injection of freshly accelerated particles.
@@
Relativistic jets launched from blazars — active galactic nuclei (AGN) with their jets oriented towards our line of sight — are known to be extremely efficient and natural particle accelerators, which are able to emit in all the electromagnetic spectrum. In fact, they are one of the few sources that can emit up to TeV energies. Blazars are also highly variable on all timescales. Depending on the properties of their optical spectrum (absence/presence of broad emission lines, see Stickel et al., 1991), they can be classified into two main sub-classes: BL Lac type objects and flat spectrum radio quasars (FSRQs) (Urry & Padovani, 1995).
@@
The broadband spectral energy distribution (SED) of blazars, extending from radio to -ray energies, has a characteristic double-peak shape. The low-energy component is produced by synchrotron radiation (see Konigl, 1981). The invoked mechanisms behind the high-energy component fall in two groups. An interpretation based on leptonic processes explains this emission through Inverse Compton (IC) scattering with low-energy photons, either synchrotron radiation via synchrotron self Compton (SSC, see Konigl, 1981; Maraschi et al., 1992) scattering, or with radiation from outside the jet via External Compton (EC, see Dermer & Schlickeiser, 1994) scattering. However, interpretations based on a hadronic nature are also used in some cases, relying on processes like proton synchrotron radiation or photo-pion production to explain the high-energy emission of blazars (see e.g. Mannheim & Biermann, 1992; Aharonian, 2000).
@@
BL Lacertae (hereafter BL Lac), the archetype of the BL Lac blazar subclass located at redshift (Miller et al., 1978), has historically shown remarkable variability with several episodes of minute- to hour-scale variations (see for instance Abeysekara et al., 2018; MAGIC Collaboration et al., 2019; Jorstad et al., 2022), becoming one of the most monitored and studied blazars over the years. It is classified as an intermediate-synchrotron-peaked (ISP) blazar according to the peak frequency of its synchrotron emission (). However, it is known for transitioning between the low-synchrotron-peaked (LSP, ) and ISP types, depending on its activity (see e.g. Ackermann et al., 2011; Nilsson et al., 2018). It has been detected several times in the very-high-energy (VHE, GeV) -ray band, always during multi-wavelength flaring emission states (Albert et al., 2007; Arlen et al., 2013). These periods of variability on very short timescales imply very compact emitting regions involving the most energetic particle acceleration processes, and challenging in some cases the theoretical models used for interpreting the broadband emission of blazars (Henri & Saugé, 2006).
@@
Here we report the results of an extreme flare observed from BL Lac by the Large-Sized Telescope prototype (LST-1), taking place between September and November 2022, and triggered after the detection of enhanced activity by the Fermi satellite (La Mura, 2022). This flare is one of the three brightest events ever detected in VHEs from BL Lac, together with similar events occurring in 2021 and 2024, and it is also the most long-lasting flaring period ever recorded for this source in the VHE domain, lasting for about 3 months, with significant variations in day, hour and minute timescales. This remarkable flare was also covered by several multi-wavelength facilities, providing a rich dataset for interpreting the multi-wavelength emission and variability of BL Lac in this period (see Sect. 2). In this work, we evaluate and interpret the VHE -ray and broadband evolution of the flare. We characterise the minute-scale -ray variability observed during this period (Sect. 3), that sets important constrains to the nature of the emitting region. The spectral behaviour of the source is evaluated (Sect. 4), and its broadband emission is modelled within AGN radiative models (Sect. 5). The main results are discussed and summarised in Sects. 6 and 7.
@@ 2.1 VHE -ray data: LST-1
LST-1 is the first of the four Large-Sized Telescopes that will be constructed in the northern site of the Cherenkov Telescope Array Observatory (CTAO, Acharyya et al., 2019) at the Roque de los Muchachos Observatory in the Canary Island of La Palma (Spain). LST-1 is currently performing commissioning and scientific observations, after the completion of its construction in October 2018. Its large collection area (23 m diameter mirror) and very sensitive camera — composed of 1855 high quantum efficiency photomultiplier tubes — allow to detect faint Cherenkov light flashes from -ray showers down to energies 20 GeV (see Abe et al., 2023).
@@
LST-1 observed BL Lac for 23 nights between September 21 and November 25, 2022 in wobble observing mode (Fomin et al., 1994), providing 36 hours of data after applying standard data quality cuts in a zenith range between 10∘ and 60∘. The analysis of the raw LST-1 data was performed with version v0.10.18 of the cta-lstchain software (López-Coto et al., 2022; Moralejo et al., 2025), including data cleaning, image parametrisation, -hadron separation and direction and energy reconstruction of the recorded events. Among this dataset, 29 hours of data were taken under dark sky conditions (no increased background light contamination due to the presence of the moon) with a standard tail-cut cleaning with thresholds of 8 and 4 photoelectrons (p.e.) for the core and neighbour pixels of the Cherenkov shower image and an intensity cut of 50 p.e. (Abe et al., 2023). The remaining 7 hours of observation were performed with different moon brightness levels. These data were analysed with cleaning thresholds ranging from 10 to 16 p.e. and 5 to 8 p.e. for the core and neighbour pixels, and increased intensity cuts between 63 and 194 p.e. due to the noise introduced by the moon, extracted from the data at each night sky background (NSB) level. The event classification was performed with a Random Forest trained using Monte Carlo (MC) simulated proton- and -like events at a declination of 34.74∘, close to that of BL Lac (42.28∘). The MCs were produced to match the NSB level of our data, accounting for the additional noise with respect to dark observations. For the event selection, energy-dependent gammaness (gamma-likeness score) and (angular distance of the detected event with respect to the nominal source position) cuts were used, optimised to preserve 70% of the simulated MC -ray events.
@@
The high level analysis was done with version 1.2 of Gammapy (Donath et al., 2023; Aguasca-Cabot et al., 2023; Acero et al., 2024). The background was characterised using an OFF region with a wobble angle of 0.4∘ and at a reflected position of 180∘ from that of BL Lac. A limit on the effective area ¿5% of the maximum area was also applied. The statistical significance was calculated following Eq. (17) from Li & Ma (1983). The energy threshold of the light curve, calculated from the maximum of the distribution of expected events from the MCs, weighted with BL Lac’s spectrum, with the highest NSB level and zenith distance between 40∘ and 50∘ to match that of the data, was estimated to be 175 GeV. Therefore, we have used a minimum energy of 200 GeV for the light curve reconstruction. The energy threshold was further verified on data against background asymmetries in the field of view (FoV). We observed a non-uniformity of the event distribution, in particular along the altitude axis. This effect increases when the telescope pointing is further from the zenith, and becomes clearly visible at zenith distances ¿40∘. Looking at the energy distribution of events in the source region and the OFF region, a clear difference in energy threshold is visible for some observations. While the difference seems mostly at energies below our analysis threshold, we also perform our spectral analysis using backgrounds counts estimation using the package BAccMod111https://github.com/mdebony/BAccMod. This package allows to extract the background distribution using the full FoV after removing the source region and can thus account for asymmetries. As expected, we observe differences in the extracted spectra below 200 GeV for some time periods, but the spectra agree above 200 GeV thus validating our results above this energy. The VHE -ray light curve is shown in the top panel of Fig. 1. We highlight October 20 and November 13 as two remarkably bright nights, when significances of 49 and 30 were reached in only 2.5 and 1.5 hours of observation. The minute-scale variability of these two nights is evaluated in Sect. 3. In particular for October 20, consisting only of dark-sky data, we adopted a threshold of 100 GeV obtained from the dark-sky MCs with zenith 20-40∘. The spectrum and SED were characterised with a log-parabola model plus a term accounting for the Extragalactic Background Light (EBL) absorption with the model from Saldana-Lopez et al. (2021). The spectral analysis is presented in detail in Sect. 4.
2.2 HE -ray data: Fermi-LAT
@@
We also analysed high-energy (HE, MeV) data taken by the pair-conversion Large Area Telescope (LAT) on board the Fermi satellite. The LAT operates in sky-survey mode in an energy range between 20 MeV and 1 TeV, with its best sensitivity around a few GeV (Atwood et al., 2009). We considered a 20∘-radius region of interest (ROI) centred at the position of BL Lac. All data from September 21 to November 30 were downloaded, selecting the Pass8 P8R3_SOURCE_V3 events between 100 MeV and 300 GeV with the Fermitools standard software. A zenith angle cut of 90∘ was applied to minimise the contamination from the limb of the Earth. Furthermore, we used the recommended versions of the Galactic diffuse emission model and isotropic component for the event selection considered222https://fermi.gsfc.nasa.gov/ssc/data/access/lat/BackgroundModels.html.
@@
To account for the emission of sources nearby BL Lac, we derived a model including all sources within the ROI, plus all sources located in an additional annular region of radius 10∘. This model was constructed using 1 year of data and a binned likelihood analysis procedure, using as reference the 4FGL-DR2 catalogue (Abdollahi et al., 2020; Ballet et al., 2020). The spectral parameters (index and normalisation) of all variable sources — as defined in the 4FGL catalogue based on the LAT variability index — contained in the ROI and with a Test Statistics (TS, Mattox et al., 1996) ¿25 (5 significance) were left as free parameters, while those not fulfilling these conditions were fixed to the catalogue values. The normalisation of the diffuse components was also left free. Convergence to the final model was achieved after removing all sources with TS¡4 (¡2).
@@
The HE -ray light curve was then obtained analysing the data within the desired time window through an unbinned likelihood analysis of events contained in 12-hour bins. For each bin we have assumed a power law spectral shape for BL Lac, leaving the spectral index and the normalisation free. The spectral parameters of the rest of the sources were fixed to the values from the model, with the exception of two highly variable and bright sources with a significant detection (TS¿25) located at a distance ¡15∘ from BL Lac (4FGL J2244.2+4057 and 4FGL J2311.0+3425). The HE -ray light curve is shown in Fig. 1.
@@ 2.3 X-ray data: Swift-XRT
The scientific payload onboard the Neil Gehrels Swift observatory includes by the X-Ray Telescope (XRT, Burrows et al., 2005), and the Optical/UV Telescope (UVOT, Roming et al., 2005). In standard operation mode, these instruments perform simultaneous observations on the same targets. In this project, we included 50 Swift-XRT and Swift-UVOT observations on BL Lac carried out since March 2022 up to January 2023.
Swift-XRT is sensitive to X-rays in the energy range. The reduced XRT data were downloaded from the UK Swift Science Data Center (Evans et al., 2009). The spectral analysis on the reduced data was performed using the XSPEC v12.13.1 package through PyXspec v2.1.2. We analysed the XRT data in the energy range from 0.3 to 10 keV. The intrinsic source emission in this range was primarily modelled with a power-law function and, to test the presence of intrinsic spectral curvature related to the transition between the synchrotron and IC bumps, with a log-parabola model. The observed spectra were fitted with the intrinsic emission models folded with the tbabs333https://heasarc.gsfc.nasa.gov/xanadu/xspec/manual/node273.html Galactic absorption term, as implemented in XSPEC. Following Raiteri et al. (2009), in the fit procedures we fixed the neutral hydrogen column density toward BL Lac to and we employed this value to correct the X-ray fluxes and SEDs for the Galactic extinction effects. The test was employed to assess whether the log-parabola model was preferred over the power-law one, with a significance level cut. The results of the X-ray analysis are summarised in Appendix B.

### 07 Apr 2026 v3

# Very-high-energy gamma-ray intranight variability from BL Lacertae during the extreme flaring state of 2022

@@ Abstract
Context. BL Lacertae (BL Lac), the archetypal blazar of its subclass and one of the most studied blazars of recent decades, has undergone a series of major multi-wavelength outbursts since 2020, resulting in its highest recorded -ray flare to date between September and November 2022, together with those from August 2021 and October 2024.
Aims. We characterised the -ray and multi-wavelength emission and spectral energy distribution (SED) of BL Lac, as well as their evolution during the major and extended -ray and multi-wavelength flare that occurred between September and November 2022.
Methods. We evaluated the variability of the flare, focusing on the nights of October 20 and November 13, 2022 when clear intranight very-high-energy (VHE, GeV) -ray variability was observed. We modelled the -ray and broadband SEDs during periods of stable emission identified with a Bayesian block analysis and interpreted the flare’s evolution in terms of the variability in the relativistic particles and the jet’s physical parameters.
Results. During this flare, the VHE emission shows an average flux of 0.23 Crab units (C.U.) above 200 GeV and a variability amplitude of more than a factor of ten. We observe intranight flux-doubling variations as fast as 8 minutes during the nights of October 20 and November 13, 2022 with maximum fluxes of 4.4 C.U. above 100 GeV and 2.8 C.U. above 200 GeV. The spectral analysis reveals a transition of the X-ray emission from the high- to the low-energy SED peak and a shift in the -ray peak towards higher energies. We interpret the broadband emission within a leptonic two-zone model in which intranight variability is explained as magnetic reconnection in a compact region closely orientated with the line of sight, while variations in the relativistic electron distributions and the injection of freshly accelerated particles explain the weekly scale variations.
@@
Relativistic jets launched from blazars, which are active galactic nuclei (AGNs) with their jets orientated towards our line of sight, are extremely efficient and natural particle accelerators that emit across the entire the electromagnetic spectrum. In fact, they are among the few sources that can emit up to TeV energies. Blazars are also highly variable on all timescales. Depending on the properties of their optical spectrum (absence or presence of broad emission lines; see Stickel et al., 1991), they can be classified into two main sub-classes: BL Lac type objects and flat spectrum radio quasars (FSRQs) (Urry and Padovani, 1995).
@@
The broadband spectral energy distribution (SED) of blazars, extending from radio to -ray energies, has a characteristic double-peak shape. The low-energy component is produced by synchrotron radiation (see Konigl, 1981). The mechanisms invoked to explain the high-energy component fall into two groups. An interpretation based on leptonic processes explains this emission through inverse Compton (IC) scattering with low-energy photons, either from synchrotron radiation via synchrotron self Compton (SSC; see Konigl, 1981; Maraschi et al., 1992) scattering or from radiation from outside the jet via external Compton (EC; see Dermer and Schlickeiser, 1994) scattering. However, hadronic-based interpretations are also used in some cases and rely on processes such as proton synchrotron radiation or photo-pion production to explain the high-energy emission of blazars (see, e.g. Mannheim and Biermann, 1992; Aharonian, 2000).
@@
BL Lacertae (hereafter BL Lac), the archetype of the BL Lac blazar subclass located at redshift (Miller et al., 1978), historically shows remarkable variability with several episodes of minute- to hour-scale variations (see, for instance Abeysekara et al., 2018; MAGIC Collaboration et al., 2019; Jorstad et al., 2022), making it one of the most monitored and studied blazars to date. It is classified as an intermediate-synchrotron-peaked (ISP) blazar according to the peak frequency of its synchrotron emission (). However, it is known to transition between low-synchrotron-peaked (LSP; ) and ISP states, depending on its activity (see, e.g. Ackermann et al., 2011; Nilsson et al., 2018). It has been detected several times in the very-high-energy (VHE; GeV) -ray band, always during multi-wavelength flaring emission states (Albert et al., 2007; Arlen et al., 2013). These periods of variability on very short timescales imply very compact emitting regions involving the most energetic particle acceleration processes, which in some cases challenge the theoretical models used to interpret the broadband emission of blazars (Henri and Saugé, 2006).
@@
Here we report the results of an extreme flare observed from BL Lac by the Large-Sized Telescope prototype (LST-1), which took place between September and November 2022 and was triggered after the detection of enhanced activity by the Fermi satellite (La Mura, 2022). This flare is one of the three brightest events ever detected in VHEs from BL Lac, together with similar events that occurred in 2021 and 2024. It is also the longest flaring period ever recorded for this source in the VHE domain, lasting about three months, with significant variations on day-, hour-, and minute-timescales. Several multi-wavelength facilities also covered this remarkable flare, providing a rich dataset for interpreting the multi-wavelength emission and variability of BL Lac during this period (see Sect. 2). In this work, we evaluate and interpret the VHE -ray and broadband evolution of the flare. We characterise the minute-scale -ray variability observed during this period (Sect. 3), which sets important constraints on the nature of the emitting region. We evaluate the spectral behaviour of the source (Sect. 4), and model its broadband emission within AGN radiative models (Sect. 5). We discuss and summarise the main results in Sects. 6 and 7.
@@ 2.1 Very-high-energy -ray data: LST-1
The LST-1 is the first of the four Large-Sized Telescopes to be constructed at the northern site of the Cherenkov Telescope Array Observatory (CTAO; Acharyya et al., 2019), located at the Roque de los Muchachos Observatory on the Canary Island of La Palma (Spain). The LST-1 is currently performing commissioning and scientific observations, following the completion of its construction in October 2018. Its large collection area (23 m diameter mirror) and highly sensitive camera, composed of 1855 high quantum efficiency photomultiplier tubes, allows detection of faint Cherenkov light flashes from -ray showers down to energies 20 GeV (see Abe et al., 2023).
@@
The LST-1 observed BL Lac for 23 nights between 21 September and 25 November 2022 in wobble observing mode (Fomin et al., 1994), providing 36 hours of data after standard data quality cuts in a zenith range between 10∘ and 60∘. We analysed raw LST-1 data with version v0.10.18 of the cta-lstchain software (López-Coto and others, 2022; Moralejo et al., 2025), including data cleaning, image parametrisation, -hadron separation, and reconstruction of the direction and energy of the recorded events. Among this dataset, 29 hours were taken under dark-sky conditions (no increased background light contamination from the Moon), using standard tail-cut cleaning with thresholds of 8 and 4 photoelectrons (p.e.) for the core and neighbour pixels of the Cherenkov shower image and an intensity cut of 50 p.e. (Abe et al., 2023). The remaining 7 hours of observation covered different moon brightness levels. These data were analysed with cleaning thresholds ranging from 10 to 16 p.e. for core pixels and 5 to 8 p.e. for neighbour pixels, and with increased intensity cuts between 63 and 194 p.e. due to noise introduced by the Moon, extracted for each night-sky background (NSB) level. The event classification used a Random Forest trained using Monte Carlo (MC) simulations of proton- and -like events at a declination of 34.74∘, close to that of BL Lac (42.28∘). We produced MC simulations to match the NSB level of our data, accounting for additional noise relative to dark observations. For event selection, we applied energy-dependent gammaness (gamma-likeness score) and (angular distance of the detected event with respect to the nominal source position) cuts, optimised to retain 70% of the simulated MC -ray events.
@@
We performed the high-level analysis with version 1.2 of Gammapy (Donath et al., 2023; Aguasca-Cabot et al., 2023; Acero et al., 2024). We characterised the background using an OFF region with a wobble angle of 0.4∘ and at a reflected position of 180∘ from BL Lac. We also applied a limit on the effective area ¿5% of the maximum area. We calculated the statistical significance following Eq. (17) from Li and Ma (1983). We estimated the light-curve energy threshold to be 175 GeV, calculated from the maximum of the expected event distribution from the MC simulations and weighted with BL Lac’s spectrum for the highest NSB level and zenith distance of ∘ to match the data. Therefore, we used a minimum energy of 200 GeV for light curve reconstruction. We further verified the energy threshold on the data against background asymmetries in the field of view (FoV). We observed a non-uniformity in the event distribution, particularly along the altitude axis. This effect increases as the telescope pointing is further from the zenith and becomes clearly visible at zenith distances ¿40∘. Examining the energy distribution of events in the source region and the OFF regions, we observe clear differences in the energy threshold for some observations. Although these differences occur mainly at energies below our analysis threshold, we also performed the spectral analysis using background counts estimated with the BAccMod package111https://github.com/mdebony/BAccMod. This package extracts the background distribution using the full FoV after removing the source region and can thus account for asymmetries. As expected, we observe differences in the extracted spectra below 200 GeV for some time periods, but the spectra agree above 200 GeV, validating our results at higher energies. We show the VHE -ray light curve in the top panel of Fig. 1. We highlight October 20 and November 13, 2022 as two remarkably bright nights, when significances of 49 and 30 were reached in only 2.5 and 1.5 hours of observation, respectively. We evaluated the minute-scale variability of these two nights in Sect. 3. For October 20, consisting only of dark-sky data, we adopted a threshold of 100 GeV, obtained from the dark-sky MC simulations at zenith angles 20-40∘. We characterised the spectrum and SED with a log-parabola model, including a term for extragalactic background light (EBL) absorption using the model from Saldana-Lopez et al. (2021). We present the spectral analysis in detail in Sect. 4.
2.2 High-energy -ray data: Fermi-LAT
@@
We analysed high-energy (HE, MeV) data taken by the pair-conversion Large Area Telescope (LAT) on board the Fermi satellite. The LAT operates in sky-survey mode over an energy range of 20 MeV and 1 TeV, with its best sensitivity around a few GeV (Atwood et al., 2009). We defined a 20∘-radius region of interest (ROI) centred on BL Lac. We downloaded all data from 21 September to 30 November, selecting the Pass8 P8R3_SOURCE_V3 events between 100 MeV and 300 GeV using the Fermitools standard software. We applied a zenith angle cut of 90∘ to minimise contamination from the Earth’s limb. We also applied the recommended versions of the Galactic diffuse emission model and the isotropic component for event selection222https://fermi.gsfc.nasa.gov/ssc/data/access/lat/BackgroundModels.html.
@@
To account for emission from sources near BL Lac, we derived a model that includes all sources within the ROI, plus all sources located in an additional annular region of radius 10∘. We constructed this model using one year of data and a binned likelihood analysis, using the second release of the Fermi-LAT Fourth Source Catalogue (4FGL-DR2) as a reference (Abdollahi et al., 2020; Ballet et al., 2020). We left the spectral parameters (index and normalisation) of all variable sources within the ROI and with a test statistics (TS; Mattox et al., 1996) ¿25 (5 significance) as free parameters, based on the 4FGL catalogue of LAT variability. We fixed the sources’ parameters not fulfilling these conditions to the catalogue values. We also left the normalisation of the diffuse components free. The final model converged after removing all sources with TS¡4 (¡2).
@@
We obtained the HE -ray light curve by analysing the data within the chosen time window using an unbinned likelihood analysis of events in 12-hour bins. For each bin, we assumed a power-law spectral shape for BL Lac, leaving the spectral index and normalisation free. We fixed the spectral parameters of the other sources to the model value, except for two highly variable and bright sources with significant detections (TS¿25) located within ¡15∘ of BL Lac (4FGL J2244.2+4057 and 4FGL J2311.0+3425). Fig. 1 shows the HE -ray light curve.
@@ 2.3 X-ray data: Swift-XRT
The scientific payload on board the Neil Gehrels Swift observatory includes the X-Ray Telescope (XRT; Burrows and others, 2005), and the Optical/UV Telescope (UVOT; Roming and others, 2005). In standard operation mode, these instruments perform simultaneous observations on the same targets. In this project, we included 50 Swift-XRT and Swift-UVOT observations of BL Lac, carried out between March 2022 and January 2023.
Swift-XRT is sensitive to X-rays in the energy range. We downloaded the reduced XRT data from the UK Swift Science Data Centre (Evans et al., 2009). We performed the spectral analysis of the reduced data using XSPEC v12.13.1 package through PyXspec v2.1.2. We analysed the XRT data in the 0.3 to 10 keV energy range. We primarily modelled the intrinsic source emission in this range with a power-law function, and to test for intrinsic spectral curvature associated with the transition between the synchrotron and IC bumps, we also employed a log-parabola model. We fitted the observed spectra with the intrinsic emission models folded with the tbabs333https://heasarc.gsfc.nasa.gov/xanadu/xspec/manual/node273.html Galactic absorption term, as implemented in XSPEC. Following Raiteri et al. (2009), we fixed the neutral hydrogen column density toward BL Lac at and used this value to correct the X-ray fluxes and SEDs for Galactic extinction effects. We employed the test to assess whether the log-parabola model is preferred over the power-law model, using a significance level threshold. We summarise the results of the X-ray analysis in Appendix B.